OpenStudy (anonymous):

Medal! Part of Part A is solved. A system of equations is shown below: -3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

Solved part of Part A: -3x + 7y = -16 -3 * 2 = -6 7 * 2 = 14 -16 * 2 = -32 New system of equations- -6x + 14y = -32 -9x + 5y = 16

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@amistre64 @Zarkon @e.mccormick @Preetha

OpenStudy (anonymous):

@dan815